OpenStudy (anonymous):

solve for x 3/x+1/2 = 5/6 6 3 9 12

OpenStudy (anonymous):

you mean this? \[x \frac{ 3 }{ x } + \frac{ 1 }{ 2 } = \frac{ 5 }{ 6 }\] ?

OpenStudy (yttrium):

Multiply the whole equation by 6x.

OpenStudy (anonymous):

yes i mean that